# ZK-APEX: Zero-Knowledge Approximate Personalized Unlearning with EXecutable Proofs

### v1

@@ 4 Proposed Approximate Unlearning Algorithm
While pruning yields a binary mask that removes parameters to induce forgetting, directly applying such a mask to a personalized model is problematic. Personalized models have adapted their parameters to sensitive local data; thus, indiscriminate pruning based on global importance scores can erase features essential for the personalization task, leading to significant accuracy degradation. This degradation arises because pruning typically targets neurons influential to the forget-set without accounting for their contribution to the retained personalized data, as neural representations often entangle multiple data sources. To mitigate this effect, we leverage the Optimal Brain Surgeon (OBS) framework lecun1989optimal; hassibi1993optimal; kuznedelev2023cap, which uses second-order information to optimally adjust the remaining weights after pruning.
@@
The final parameters $\theta_{u}=\theta_{p}+\delta w^{\star}$ obey $(\theta_{u})_{M}=0$. In practice, we compute $C_{p}^{-1}E_{M}$ via damped Fisher-vector products and conjugate gradients, and invert only the small $|M|\times|M|$ Schur complement. Following prior work kurtic2022optimal; kuznedelev2023cap, we adopt a block-wise Fisher structure for computational efficiency while accurately capturing the user-specific curvature.
@@ 6.2 Baselines
Approximate unlearning. We evaluate two gradient-based baselines. (i) Gradient Ascent (GA) graves2021amnesiac; thudi2022unrolling increases the loss on the forget-set without accounting for retained data. (ii) SCRUB kurmanji2023towards alternates ascent on the forget-set and descent on the retain set to balance forgetting and utility. Both require iterative optimization, making them expensive to verify in zero-knowledge circuits.
Exact unlearning. As a reference, we retrain the model on the retain set and then re-personalize on each client, following eisenhofer2025verifiable. This achieves exact unlearning but is computationally infeasible for edge deployment.
@@ 6.3 Experimental Setup
Models and datasets.
We evaluate two models: Google’s ViT-B/16 for classification and Meta’s OPT-125 LLM. Each model was fine-tuned on a separate personalization set to emulate the provider–client setup. ViT, pretrained on ImageNet, was personalized on ImageNet-Sketch, while OPT-125 was fine-tuned on Scala, C, C++, and Java before being personalized on Rust using CodeParrot’s GitHub Clean dataset. The initial fine-tuning narrows OPT-125’s broad training scope, enabling a distinct forget set. Both setups capture domain shifts typical of client-side personalization.
Unlearning and compensation.
A forget-set, $D_{f}$, consisting of 33,600 samples, or 2.6% of the original data, was selected for the ViT. For the LLM, $D_{f}$ was chosen to be a subset of the Scala training examples, representing 4.8% of the original Scala train set and 1.2% of the overall train set.
In both cases an unlearning mask is computed on the pretrained model, targeting only the MLP sublayers across all transformer blocks and pruning $4\%$ of their parameters while leaving attention heads intact, following pochinkov2024dissecting which shows MLP pruning to be more effective for unlearning. The saliency score in Section 4 is used with curvature estimated from a diagonal block-wise empirical Fisher All hyperparameters are tuned on a separate validation split, and sensitivity analyses are reported in Section G
Evaluation.
For the image classification task, accuracy was used as the primary metric to evaluate performance. Membership inference (MIA) kurmanji2023towards leakage was also evaluated For the generation task, both Top-1 accuracy and perplexity are used.
Table 1: Main results. Left: unlearning efficacy (Forget = forget-set accuracy, Pers. = personalization accuracy, AUC = membership inference AUC). Right: ZKP efficiency (Time = proving time in hours, Mem = peak prover memory in GB, Size = proof size in MB, Verify = verification time in minutes).
Method Unlearning ZK Proof Efficiency
Forget (%) $\downarrow$ Pers. (%) $\uparrow$ AUC $\downarrow$ Time (h) $\downarrow$ Mem (GB) $\downarrow$ Size (MB) $\downarrow$ Verify (min) $\downarrow$
@@
Table 2: Unlearning efficacy on LLM. Comparison of the forgetting–personalization trade-off.
Method Unlearning Efficacy
Forget Acc. (%) $\downarrow$ Forget PPL $\uparrow$ Personal Acc. (%) $\uparrow$ Personal PPL $\downarrow$
@@ 6.4 Main Results
ZK proof generation efficiency. Table 1 summarizes the computational cost of generating and verifying ZKP for different unlearning operators. For approximate (GA, SCRUB) and exact unlearning baselines, the gradient computation circuits exceeded available memory; following standard practice zcash-halo2; maheri2025telesparse, we partitioned the computation into multiple sub circuits, each capped at $2^{20}$ rows to fit within 256,GB RAM. The per-sample proving time, proof size, and verification time were measured for each sub-circuit and then scaled by the number of sub-circuits per sample and by the total number of samples in each unlearning or learning procedure.
Our method achieves orders-of-magnitude lower proving time and memory usage compared to optimization-based baselines. This improvement arises from its linear operation formulation, which avoids iterative updates or gradient reconstruction inside the circuit. Overall, the results confirm that efficient and verifiable personalized unlearning is feasible on high-capacity models without compromising performance or proof succinctness.
@@ 6.5 Edge-Device Evaluation
Table 3: Edge-device proof generation overhead. All values correspond to a single ViT-B/16 block update
Fisher Block Size Sparsity Ratio ($k$) Proving Time (h) Peak Memory (MB) Proof Size (KB)
@@
Despite the computational limits of mobile hardware, proof generation remains practical. As shown in Table 3, proving time scales nearly linearly with the pruning ratio $k$ and sub-quadratically with the Fisher block size, while memory and proof sizes stay within device capacity. For a ViT-B/16 block, end-to-end proof generation completes within a few hours without requiring high memory and can further reduce wall-clock time by parallelizing block proofs. Because ZK-SNARK verification is lightweight, the provider can validate proofs from many clients in parallel. These results confirm that the proposed ZK-APEX framework is feasible for edge intelligence deployments, such as on-device personalization in transformer based Apple Intelligence gunter2024apple, where users can locally prove correct unlearning while the provider verifies multiple clients concurrently.
@@ 7 Conclusion and Future Work
This work introduced the first framework for verifiable personalized unlearning on edge devices, motivated by the need to enforce user data deletion under privacy and regulatory constraints without relying on trust in local computation. We proposed a pruning-based approximate unlearning algorithm with OBS compensation, designed to be zero-shot (requiring no retraining iterations) and inherently compatible with ZKP systems. The approach enables providers to verify, in zero knowledge, that a client correctly executed an agreed unlearning transformation on their personalized model without revealing weights or data. Our linear operation formulation avoids stochastic optimization inside the circuit, eliminating vulnerabilities linked to randomness in SGD and thereby improving robustness against forging attacks on verifiable unlearning.
@@
Empirical results on personalized ViT models fine-tuned on ImageNet-Sketch confirm that the proposed method effectively removes forget-set influence while recovering over 99% of the personalized accuracy lost through naïve pruning. These findings demonstrate that efficient, privacy-preserving verification of approximate unlearning is feasible even for large transformer architectures on mobile phone.
@@
Our formulation assumes that residual-gradient, cross-curvature, and quadratic effects on unmasked parameters are minor, preventing compensation from negating forgetting. An extension is to project compensation onto the orthogonal subspace of forget directions, canceling these terms while staying computationally light and ZK-compatible. Exploring alternative proving systems such as MPC- and polynomial-commitment-based SNARKs could further optimize latency and scalability. From a privacy perspective, adding differential privacy to the unlearning mask can defend against inversion and reconstruction attacks, mitigating verification-forging risks DBLP:conf/icml/ZhangCSL24. At the system level, hardware-assisted secure erasure complements our cryptographic verification by guaranteeing complete data removal. Finally, when personalization and forget-set distributions overlap, balancing retention and forgetting becomes difficult. Determining whether approximate unlearning suffices or exact retraining remains necessary in such regimes is an open question for future work.
@@ Appendix B Background: Circuits and Zero-Knowledge Proof Systems
This holds if and only if the witness satisfies the entire circuit, thereby transforming constraint satisfaction into a single polynomial divisibility condition groth2016size.
@@
Halo2 and modern polynomial commitments.  Halo2 zcash-halo2 extends the above paradigm using transparent polynomial commitment schemes (e.g., KZG-type) that avoid trusted setup and support recursive proofs. It represents circuits as constraint systems over field polynomials and verifies each constraint through low-degree testing. Halo2 thus achieves scalability while preserving soundness, correctness, and zero-knowledge under standard cryptographic assumptions. These three properties, informally ensure that proofs are both reliable and privacy-preserving:

### v2

@@ 4 Proposed Approximate Unlearning Algorithm
While pruning yields a binary mask that removes parameters to induce forgetting, directly applying such a mask to a personalized model is problematic. Personalized models have adapted their parameters to sensitive local data; thus, indiscriminate pruning based on global importance scores can erase features essential for the personalization task, leading to significant accuracy degradation. This degradation arises because pruning typically targets neurons influential to the forget-set without accounting for their contribution to the retained personalized data, as neural representations often entangle multiple data sources. To mitigate this effect, we leverage the Optimal Brain Surgeon (OBS) framework LeCun et al. (1989); Hassibi et al. (1993); Kuznedelev et al. (2023), which uses second-order information to optimally adjust the remaining weights after pruning.
@@
The final parameters $\theta_{u}=\theta_{p}+\delta w^{\star}$ obey $(\theta_{u})_{M}=0$. In practice, we compute $C_{p}^{-1}E_{M}$ via damped Fisher-vector products and conjugate gradients, and invert only the small $|M|\times|M|$ Schur complement. Following prior work Kurtic et al. (2022); Kuznedelev et al. (2023), we adopt a block-wise Fisher structure for computational efficiency while accurately capturing the user-specific curvature.
@@ 6.2 Baselines
Approximate unlearning. We evaluate two gradient-based baselines. (i) Gradient Ascent (GA) Graves et al. (2021); Thudi et al. (2022a) increases the loss on the forget-set without accounting for retained data. (ii) SCRUB Kurmanji et al. (2023) alternates ascent on the forget-set and descent on the retain set to balance forgetting and utility. Both require iterative optimization, making them expensive to verify in zero-knowledge circuits.
Exact unlearning. As a reference, we retrain the model on the retain set and then re-personalize on each client, following Eisenhofer et al. (2025). This achieves exact unlearning but is computationally infeasible for edge deployment.
@@ 6.3 Experimental Setup
Table 1: Unlearning efficacy and ZK proof efficiency. Left (unlearning): lower is better for Forget Acc. and MIA AUC; higher is better for Personal Acc. (all in %). Right (ZKP): lower is better for proving time (h), peak memory (GB), proof size (MB), and verification time (min).
Method Unlearning Efficacy ZK Proof Efficiency
Forget Acc. Personal Acc. MIA AUC Proving Time Peak Mem. Proof Size Verification Time
@@
Models and datasets.
We evaluate two models: Google’s ViT-B/16 for classification and Meta’s OPT-125M for language modeling. Each model was fine-tuned on a separate personalization set to emulate the provider–client setup. ViT, pretrained on ImageNet, was personalized on ImageNet-Sketch using full fine-tuning; in Appendix I, we additionally evaluate AdaptFormer Chen et al. (2022) to demonstrate the extensibility of our method across different personalization algorithms. For OPT-125M, we first fine-tune on Scala, C, C++, and Java using LoRA Hu et al. (2022) to narrow its broad pretraining scope and enable a distinct forget-set, and then personalize on Rust from CodeParrot’s GitHub Clean dataset, again using LoRA Hu et al. (2022). Both setups capture domain shifts typical of client-side personalization.
Unlearning and compensation.
A forget-set, $D_{f}$, consisting of 33,600 samples (2.6% of the original training data) was selected for the ViT. For the LLM, $D_{f}$ was chosen as a subset of the Scala training examples, representing 4.8% of the Scala training split and 1.2% of the overall training data. Unless stated otherwise, we construct $D_{f}$ by uniform random sampling from the designated source split. To study a structured (class-conditional) forgetting regime, we additionally evaluate cases where $D_{f}$ is concentrated in a single (or a small number of) classes; see Appendix J.
In both cases an unlearning mask is computed on the pretrained model, targeting only the MLP sublayers across all transformer blocks and pruning $4\%$ of their parameters while leaving attention heads intact, following Pochinkov and Schoots (2024); Meng et al. (2022) which show MLP pruning to be more effective for unlearning. The saliency score in Section 4 is used with curvature estimated from a diagonal block-wise empirical Fisher All hyperparameters are tuned on a separate validation split, and sensitivity analyses are reported in Section G
Appendix K further studies sensitivity to the forget-set size $|D_{f}|$ by increasing $|D_{f}|$ beyond our default setting and tracing the resulting forgetting–personalization accuracy trade-off. In particular, we report how forgetting effectiveness and retention effect as $|D_{f}|$ grows, providing a clearer picture of robustness under larger unlearning requests.
Evaluation.
For the image-classification task, we use Top-1 accuracy as the primary utility metric. We additionally evaluate black-box membership leakage via a membership inference attack (MIA) adapted to unlearning Kurmanji et al. (2023): we form an attack dataset using the unlearned model’s per-example output loss on forget-set samples (members) versus an identically distributed held-out test set (non-members), train an attack classifier on a class-balanced split, and report its performance on a disjoint held-out evaluation split using the area under the receiver operating characteristic curve (AUC) (values closer to $50\%$ indicate a weaker membership signal). This diagnostic measures distinguishability from model outputs and is independent of the zero-knowledge privacy of the proof transcript. For the generation task, we report both Top-1 accuracy and perplexity (PPL).
Table 2: Unlearning efficacy on LLM. Forgetting vs. personalization trade-off. Acc. is in % (higher is better); PPL is perplexity (lower is better).
Method Forget Personal
Acc. PPL Acc. PPL
@@ 6.4 Main Results
ZK proof generation efficiency. Table 1 summarizes the computational cost of generating and verifying ZKP for different unlearning operators. In our deployment setting, the reported costs correspond to an unlearning window/event: the provider batches deletion requests within a window, publishes a single public artifact $\Psi$ for that window, and each client generates one proof certifying $\theta_{u}=U(\theta_{p};\Psi)$ for its locally personalized model (rather than producing a separate proof per individual deletion request). For approximate (GA, SCRUB) and exact unlearning baselines, the gradient computation circuits exceeded available memory; following standard practice Zcash (2022); Maheri et al. (2025), we partitioned the computation into multiple sub circuits, each capped at $2^{20}$ rows to fit within 256,GB RAM. The per-sample proving time, proof size, and verification time were measured for each sub-circuit and then scaled by the number of sub-circuits per sample and by the total number of samples in each unlearning or learning procedure. A brief note on batching and the resulting throughput/communication interpretation is provided in Appendix H.
Our method achieves orders-of-magnitude lower proving time and memory usage than optimization-based baselines. This improvement stems from its linear formulation, which avoids iterative updates and gradient reconstruction inside the circuit. The results confirm that efficient, verifiable personalized unlearning is feasible for high-capacity models without compromising performance or proof succinctness.
@@ 6.5 Edge-Device Evaluation
Table 3: Edge-device proof overhead. Per-block (ViT-B/16) proving cost on the client device.
Fisher block $k$ Proving time Peak mem. Proof size
@@
Despite the computational limits of mobile hardware, proof generation remains practical. As shown in Table 3, proving time scales nearly linearly with the pruning ratio $k$ and sub-quadratically with the Fisher block size, while memory and proof sizes stay within device capacity. For a ViT-B/16 block, end-to-end proof generation completes within a few hours without requiring high memory and can further reduce wall-clock time by parallelizing block proofs. Because ZK-SNARK verification is lightweight, the provider can validate proofs from many clients in parallel. These results confirm that the proposed ZK-APEX framework is feasible for edge intelligence deployments, such as on-device personalization in transformer based Apple Intelligence Gunter et al. (2024), where users can locally prove correct unlearning while the provider verifies multiple clients concurrently.
@@ 7 Conclusion and Future Work
This work introduced the first framework for verifiable personalized unlearning on edge devices, motivated by the need to enforce user-data deletion under privacy and regulatory constraints without relying on trust in local computation. We proposed a pruning-based approximate unlearning algorithm with OBS compensation, designed to be zero-shot (requiring no retraining iterations) and inherently compatible with ZKP systems. The approach enables providers to verify, in zero knowledge, that a client correctly executed an agreed unlearning transformation on its personalized model without revealing model weights or private data. Our linear operation formulation avoids stochastic optimization inside the circuit, eliminating vulnerabilities linked to randomness in SGD and thereby improving robustness against forging attacks on verifiable unlearning.
@@
Empirical results on personalized ViT models fine-tuned on ImageNet-Sketch confirm that the proposed method effectively removes forget-set influence while recovering over $99\%$ of the personalized accuracy lost through naive pruning. These findings demonstrate that efficient, privacy-preserving verification of approximate unlearning is feasible even for large transformer architectures on mobile devices.
@@
Several directions remain for future research. Our current formulation assumes that residual-gradient, cross-curvature, and quadratic effects on unmasked parameters are sufficiently small, so that compensation does not negate forgetting. One extension is to project compensation onto subspaces orthogonal to forget directions, canceling these terms while remaining computationally light and ZK-compatible. Exploring alternative proving systems, such as MPC-based or polynomial-commitment-based SNARKs, could further improve scalability and reduce proof latency. From a privacy perspective, incorporating differential privacy into mask construction may help defend against inversion and reconstruction attacks during verification Zhang et al. (2024); mahdi Maheri et al. . At the systems level, hardware-assisted secure erasure would complement our cryptographic verification by ensuring that prior model states are permanently removed from device storage.
Another promising direction is to extend the framework beyond sample-defined forget sets to richer forms of forgetting. In particular, concept-level unlearning aims to remove knowledge associated with semantic concepts rather than individual samples, while feature-level unlearning targets finer-grained attributes or representation components. Adapting verifiable unlearning mechanisms to support these richer request types remains an open challenge. More broadly, it would be valuable to support additional deletion strategies within the same verifiable interface, provided they can be expressed as deterministic, circuit-checkable operators. Finally, when personalization and forget-set distributions overlap substantially, balancing retention and forgetting becomes inherently difficult. Determining whether approximate unlearning suffices, or whether exact retraining remains necessary in such regimes, is an important open question for future work.
ACKNOWLEDGEMENTS
We wish to acknowledge the thorough and useful feedback from anonymous reviewers and our shepherd. The research in this paper was supported by the UKRI Open Plus Fellowship (EP/W005271/1 Securing the Next Billion Consumer Devices on the Edge) and EU CHIST-ERA GNNs for Network Security and Privacy (GRAPHS4SEC) projects. Alex Davidson’s work is funded by national funds through FCT — Fundação para a Ciência e a Tecnologia, I.P., under the project “Fully-Homomorphic Encryption from Post-Quantum Code-Based Assumptions”, ref. 2024.12595.CMU, DOI 10.54499/2024.12595.CMU, and under the LASIGE Research Unit, ref. UID/00408/2025, DOI 10.54499/2024.07643.IACDC, and the LASIGE Research Unit, ref. UID/00408/2025 – LASIGE.
@@ Appendix B Background: Circuits and Zero-Knowledge Proof Systems
This holds if and only if the witness satisfies the entire circuit, thereby transforming constraint satisfaction into a single polynomial divisibility condition Groth (2016).
@@
Halo2 and modern polynomial commitments.  Halo2 Zcash (2022) extends the above paradigm using transparent polynomial commitment schemes (e.g., KZG-type) that avoid trusted setup and support recursive proofs. It represents circuits as constraint systems over field polynomials and verifies each constraint through low-degree testing. Halo2 thus achieves scalability while preserving soundness, correctness, and zero-knowledge under standard cryptographic assumptions. These three properties, informally ensure that proofs are both reliable and privacy-preserving:
@@ Appendix F Notation Table
Appendix G Ablations and Sensitivity Study
Appendix H Deployment model and throughput/communication interpretation
We consider a windowed deployment in which deletion requests are accumulated over a fixed interval and aggregated into a forget-set $D_{f}$. For each window, the provider publishes a single public traceability artifact $\Psi$ (e.g., a sparse mask), and each client produces a proof certifying that its released update satisfies $\theta_{u}=U(\theta_{p};\Psi)$ for its locally personalized model. Accordingly, the costs in Table 1 should be interpreted at the granularity of one client proof per window.
A minimal systems interpretation follows from two measured quantities in Table 1: proof size $S$ and verification time $t_{\mathrm{ver}}$. On the client side, the upload time under uplink bandwidth $b$ (bits/s) is
$t_{\mathrm{up}}\approx\frac{8S}{b}.$
For the ViT-B/16 configuration in Table 1 ($S\approx 401$ MB), this corresponds to approximately 2233 minutes at b=20b=202525 Mbps. On the verifier side, a single verifier instance has throughput
$\lambda_{\mathrm{ver}}\approx\frac{1}{t_{\mathrm{ver}}},$
so with $t_{\mathrm{ver}}\approx 10$ minutes the instance processes roughly one proof per $\sim 10$ minutes; throughput scales linearly with the number of verifier instances.
The reported $S$ and $t_{\mathrm{ver}}$ are totals across all verified model blocks. Since our prover is block-decomposed, block proofs can be generated in parallel and streamed as they are produced, allowing overlap between proving and transmission and improving wall-clock latency in practice. If communication becomes the dominant bottleneck, recursive composition can aggregate block proofs into a single succinct proof to reduce transmitted proof material; we view this as a systems optimization direction and report unaggregated costs in Table 1.
Appendix I Beyond Full Fine-Tuning: AdaptFormer Personalization
In the main ViT experiments (Section 6.4), we instantiate client-side personalization using full fine-tuning. However, ZK-APEX is defined as a feasible, deterministic unlearning operator $U(\theta_{p};\Psi)$ acting on the resulting personalized parameters (Section 3): the provider publishes a public traceability artifact $\Psi$ (a sparse mask), and the client computes a curvature-aware compensation from $D_{p}$ (Group-OBS) and proves operator compliance in zero knowledge. Since this workflow depends on the personalized model $\theta_{p}$ but not on the specific optimization procedure used to obtain $\theta_{p}$ from $\theta_{0}$, it should extend to other widely used personalization methods. To validate this claim, we repeat the ViT pipeline while changing only the personalization operator $P$: instead of full fine-tuning, we personalize ViT-B/16 on ImageNet-Sketch using AdaptFormer Chen et al. (2022), while keeping the unlearning configuration identical to Section 6.3 (mask computed on the pretrained model, targeting MLP sublayers across Transformer blocks, with the same pruning ratio and the same Group-OBS compensation and ZK statement).
AdaptFormer Chen et al. (2022) is a parameter-efficient adaptation method for vision transformers that freezes the pretrained backbone and injects lightweight bottleneck “adapter” modules into Transformer blocks (typically as an additional residual branch alongside the block MLP/FFN), training only the adapter parameters. A typical adapter computes a down-projection to a low-dimensional bottleneck, applies a nonlinearity, and projects back to the model dimension, which is then added (with a learnable scale) to the block representation, enabling scalable adaptation with a small number of trainable parameters relative to full fine-tuning. Table 4 reports the forgetting–personalization trade-off under AdaptFormer personalization, using the same metrics as Table 1 but focusing on accuracy: “Pre-unlearning model” refers to the AdaptFormer-personalized $\theta_{p}$ prior to unlearning, “Exact unlearning” is the gold-standard retrain-on-retain then re-personalize using the same AdaptFormer operator $P$ (Equation 2), and ZK-APEX applies our masked unlearning plus Group-OBS compensation starting from the AdaptFormer-personalized model.
Table 4: Unlearning efficacy on ViT under AdaptFormer personalization. Comparison of the forgetting–personalization trade-off.
Method Unlearning Efficacy
Forget Acc. (%) $\downarrow$ Personal Acc. (%) $\uparrow$
Pre-unlearning 92.5 82.5
Mask only (4%) 54.8 77.9
ZK-APEX 54.7 81.4
Exact unlearning 54.7 82.3
Under AdaptFormer personalization (Table 4), we observe the same qualitative behavior as with full fine-tuning. Mask-only unlearning attains comparable forgetting but incurs an approximately 4.4-point drop in personalization accuracy relative to the exact unlearning gold standard. In contrast, ZK-APEX recovers about 3.5 points of this gap (roughly 80% recovery), while matching the gold standard on forgetting. Overall, this is consistent with the full fine-tuning results: ZK-APEX substantially narrows the retention gap between naive masking and retrain-and-repersonalize, supporting its applicability across different personalization algorithms.
Appendix J Structured Forget Sets: Per-Class and Few-Class Unlearning
In the main evaluation, we construct the forget-set $D_{f}$ by sampling examples uniformly at random from the designated source split (Section 6.3), which typically yields a class-diverse forget-set. Real unlearning requests can be more structured and thus more correlated, e.g., when a request targets data associated with a particular class or a small set of related labels. To probe this regime, we consider class-conditional forget sets where $D_{f}$ is restricted to fewer classes, increasing intra-forget-set correlation. Concretely, we evaluate two settings on ViT: (i) 1-class forgetting, where $D_{f}$ consists of 600 examples from a single class, and (ii) 8-class forgetting, where $D_{f}$ consists of 600 examples from each of 8 distinct classes (4,800 total). To enable comparisons across different forget-set constructions, Table 5 reports the relative deviation of ZK-APEX from the Exact unlearning gold standard, computed as $(\mathrm{metric}(\textsf{ZK\text{-}APEX})-\mathrm{metric}(\textit{Exact}))/\mathrm{metric}(\textit{Exact})$ for both forget accuracy and personalization accuracy; the Random row corresponds to the main setting (Table 1). With this normalization, negative values for forget accuracy indicate stronger forgetting than the gold standard, while positive values for personalization accuracy indicate better retention.
For these structured forget sets, we keep the provider-side masking procedure fixed (i.e., the same masking construction and selection rule as in the main experiments) and only retune the client-side compensation hyperparameters. This aligns with the provider–client workflow in our threat model: the provider publishes a single public traceability artifact $\Psi$ that is not customized per client, whereas each client can locally choose compensation settings to best preserve its personalization utility under the imposed unlearning request. This protocol isolates the impact of increased forget-set structure on the forgetting–retention trade-off without relying on per-client tuning of the masking stage.
Table 5: Structured forget sets: relative deviation from exact unlearning. Relative differences (%) between ZK-APEX and Exact unlearning across forget-set constructions.
Forget-set Relative diff. to Exact (%)
$\bm{\Delta_{\mathrm{forget}}}$ ($\downarrow$) $\bm{\Delta_{\mathrm{personal}}}$ ($\uparrow$)
Random (Table 1) + 0.20 - 0.12
8-class + 0.25 + 0.12
1-class 0.00 + 0.37
As shown in Table 5, ZK-APEX remains tightly aligned with the Exact unlearning gold-standard under structured (class-conditional) forget sets, and the alignment is slightly stronger when the forget set is concentrated in fewer classes. A plausible explanation is that increasing intra-$D_{f}$ correlation makes the forget signal more spatially and semantically localized in the representation space, so the provider mask can more cleanly isolate parameters tied to the targeted content while leaving client-relevant directions intact. Interestingly, in the 8-class and especially 1-class regimes, the masked update with compensation not only preserves personalization relative to the gold-standard but slightly improves it, consistent with the broader observation that mild sparsification/pruning can act as an implicit regularizer and occasionally improve generalization (and thus accuracy) rather than only compressing the model Bartoldson et al. (2020); Frankle and Carbin (2018).
Appendix K Sensitivity to Forget-Set Size
In the main ViT evaluation (Table 1), we consider a default forget-set size of $|D_{f}|/|D|=2.6\%$. In practice, unlearning requests may be substantially larger, and it is important to understand how the forgetting–personalization trade-off scales as a larger fraction of the training data is removed. To this end, we increase the forget-set fraction from the default setting up to $20\%$ by sampling $D_{f}$ uniformly at random from the same source split as in Section 6.3, and evaluate ZK-APEX across these regimes. For each forget-set fraction, we compare against the Exact unlearning oracle (retrain on $D\setminus D_{f}$ and then re-personalize) and report results using the same metrics as the main ViT experiments.
Figure 6: Scaling $|D_{f}|$: relative forget accuracy vs. exact unlearning. Each point corresponds to a different forget-set fraction (up to $20\%$). Lower (more negative) values indicate stronger forgetting relative to the Exact unlearning oracle.
To make comparisons across different $|D_{f}|$ values scale-free, we plot the relative deviation of ZK-APEX from Exact unlearning for both forget accuracy and personalization accuracy, using the same normalization as in Section J. Concretely, for a metric $m(\cdot)$ we compute
$\Delta_{m}(\%)\;=\;100\cdot\frac{m(\textsf{ZK\text{-}APEX})-m(\textit{Exact})}{m(\textit{Exact})}.$
We report $\Delta_{\mathrm{forget}}$ for forget accuracy and $\Delta_{\mathrm{personal}}$ for personalization accuracy. Under this convention, negative $\Delta_{\mathrm{forget}}$ indicates stronger forgetting than the oracle (lower forget accuracy), while positive $\Delta_{\mathrm{personal}}$ indicates better retention than the oracle (higher personalization accuracy). Figure 6 summarizes how these relative deviations evolve as the forget-set fraction increases up to $20\%$.
Across increasing forget-set sizes (up to $20\%$ of training data), personalization remains relatively stable: the relative deviation in personalization accuracy between ZK-APEX and Exact unlearning stays below $1\%$ in all cases, indicating that the client-side compensation continues to preserve personalized utility as $|D_{f}|$ grows. On the forgetting side, ZK-APEX closely tracks the oracle for small-to-moderate forget ratios (up to $10\%$), with only negligible relative differences. The only clear degradation appears at the largest forget ratio ($20\%$), where forgetting becomes noticeably weaker. This behavior is consistent with our protocol in which the provider-side masking hyperparameter is fixed (e.g., $k=4\%$): for very large unlearning requests, a fixed mask budget can become a bottleneck, limiting how much of the oracle’s unlearning update can be expressed, whereas for smaller requests it remains sufficient to match the oracle while maintaining retention.